# Ultrafast optical switching to a heterochiral charge-density wave state

### v1

@@
Ultrafast optical switching to a heterochiral charge-density wave state
Wayne Cheng-Wei Huang Department of Ultrafast Dynamics, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany IV. Physical Institute–Solids and Nanostructures, University of Göttingen, 37077 Göttingen, Germany Institute of Physics, Academia Sinica, Taipei 115201, Taiwan (R.O.C.)    Sai Mu Center for Experimental Nanoscale Physics, Department of Physics and Astronomy, University of South Carolina, Columbia, South Carolina 29208, USA    Gevin von Witte Department of Ultrafast Dynamics, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany Institute for Biomedical Engineering, University and ETH Zurich, Zurich 8092, Switzerland Institute of Molecular Physical Science, ETH Zurich, Zurich 8093, Switzerland    Yanshuo Sophie Li IV. Physical Institute–Solids and Nanostructures, University of Göttingen, 37077 Göttingen, Germany    Felix Kurtz Department of Ultrafast Dynamics, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany    Sheng-Hsiung Hung Department of Physics, National Tsing Hua University, Hsinchu 30013, Taiwan (R.O.C.)    Horng-Tay Jeng Department of Physics, National Tsing Hua University, Hsinchu 30013, Taiwan (R.O.C.) Institute of Physics, Academia Sinica, Taipei 115201, Taiwan (R.O.C.) Physics Division, National Center for Theoretical Sciences, Taipei 10617, Taiwan (R.O.C.)    Kai Rossnagel Institute of Experimental and Applied Physics, Kiel University, 24098 Kiel, Germany Ruprecht Haensel Laboratory, Deutsches Elektronen-Synchrotron DESY, 22607 Hamburg, Germany    Jan Gerrit Horstmann Department of Ultrafast Dynamics, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany Department of Materials, ETH Zurich, 8093 Zurich, Switzerland    Claus Ropers Email: claus.ropers@mpinat.mpg.de Department of Ultrafast Dynamics, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany IV. Physical Institute–Solids and Nanostructures, University of Göttingen, 37077 Göttingen, Germany
Optical control of correlated electronic states promises unprecedented tunability of novel functional materials Bao2022 ; Torre2021 ; Basov2017 . Tailored optical excitations can steer a system along non-equilibrium pathways to metastable states with specific structural or electronic properties Stojchevska2014 ; Horstmann2020 ; Rini2007 . A much-desired feature is the reproducible and ultrafast switching to functional states. The light-induced hidden state of 1T-TaS2, with its strongly enhanced conductivity and exceptionally long lifetime, represents a unique model system for studying the switching of correlated electronic states using femtosecond optical stimuli Stojchevska2014 ; Maklar2023 . However, despite intense investigation, the switching mechanism and the structural origins of the distinctive electronic properties of the hidden state have not been fully uncovered. Here, we use surface-sensitive electron diffraction in combination with a femtosecond optical quench to reveal coexistent charge-density wave chiralities as a new structural feature of the hidden state. We find that a single-pulse optical quench produces a state with long-range structural order and different weights of the two chiral enantiomorphs of the charge-density wave. Harnessing a double-pulse optical quench Horstmann2020 ; Maklar2023 ; Hase2015 ; Weiner1990 , we trace the origin of the mixed chirality to the transient electronic excitation of the host crystal. The coexistent long-range-order of both chiralities suggests the presence of extended heterochiral charge-density wave interfaces, which results in a higher-level, fractal-type moiré superstructure. Density functional theory simulations for such a charge-density wave moiré superstructure reveal multiple flat bands, Dirac cones, and a kagome electronic subsystem around the Fermi energy. Our findings shed light on novel electronic properties gained by chiral interface engineering, and create avenues for light-induced moiré superstructures in quasi-two-dimensional materials.
@@
Ultrafast control of functional materials via non-equilibrium pathways promises access to tailored electronic and structural properties Bao2022 ; Torre2021 ; Basov2017 . A particularly prominent example is the 1T-TaS2 hidden state Stojchevska2014 , which may enable novel non-volatile all-electronic memory devices Vaskivskyi2016 . The long-lived metastable hidden state exhibits metallic conductivity Stojchevska2014 ; Gao2022 ; Svetin2017 and is generated from the insulating low-temperature commensurate charge-density wave phase (C-phase) vonWitte2019 by applying single laser or voltage pulses Stojchevska2014 ; Gerasimenko2019 ; Cho2016 ; Ma2016 ; Vaskivskyi2015 . Structural changes associated with this state have been probed by scanning tunneling microscopy (STM) and X-ray diffraction, yielding an emergent mosaic network of domain walls and the breakdown of interlayer dimerization Gerasimenko2019 ; Cho2016 ; Ma2016 ; Stahl2020 . However, a unifying picture that explains the structural origin of metallic conductivity and long-term stability, as well as the possible role of quench-induced disorder, remains to be developed.
@@
In the present work, we reveal a new structural characteristic concurring with the hidden state: coexistent charge-density wave (CDW) chiralities. Using low-energy electron diffraction (LEED) Vogelgesang2018 following a single-pulse optical quench Stojchevska2014 ; Vaskivskyi2015 ; Gao2022 ; Gerasimenko2019 ; Stahl2020 , our experiments demonstrate the emergence of a long-range-ordered CDW chirality opposite to that of the initial state. The electronic origin of this structural transformation is verified by double-pulse excitation Maklar2023 ; Ravnik2018 . Fluence-dependent diffraction intensities of the newly formed chirality show a threshold behavior for 13 measured samples. The identified threshold fluence varies by about a factor of 3 among the ensemble of crystals, with an average value of 1.2 mJ/cm-${}^{\textrm{-}2}$, in accordance with previously observed thresholds for hidden-state generation Stojchevska2014 ; Gerasimenko2019 ; Stahl2020 . At high fluence, most specimens exhibit a saturated diffraction intensity of the minority chirality between 2 % and 7 % of that of the majority chirality, while two samples reach about 10 % without a clear sign of saturation. Our study implies extended CDW moiré interfaces present in the laser-induced state, for which DFT simulations predict a new type of kagome charge order that supports in-plane metallic conductivity. Alongside disordered CDW stacking arrangements Cho2016 ; Ma2016 , such heterochiral interfaces may contribute to the electronic properties of the hidden state, and more generally, suggest active laser control of interface chirality also in other contexts Ohta2021 ; Stojchevska2018 ; Song2022 ; Liu2023 .
@@
Figure 1: LEED images of the laser-induced heterochiral CDW state. a, Sketch of the optical excitation sequence, with quadrants representing the LEED images in d. b, Experimental setup for single- and double-pulse quench. c, The two chiralities of the 1T-TaS2 CDW superstructure, rotated by ±$\pm 13.9^{\circ}$ relative to the host lattice (black line). d, Sections of LEED images in the sequence from the C-phase to the heterochiral CDW state and back to a monochiral state. A single-pulse quench of the C-phase (top-left panel, sample #1, pulse fluence: 1.7 mJ cm-${}^{\textrm{-}2}$) generates first- and second-order $\beta$ superstructure peaks (solid and dashed circles, respectively, in the top-right panel). Annealing with a laser pulse train reverses the sample to a monochiral state (bottom-left panel) indistinguishable from the initial C-phase. A difference image (bottom-right panel) highlights the emergence of $\beta$ peaks and the suppression of $\alpha$ peaks in the heterochiral state. e, Close-up LEED images of the C-phase and the heterochiral state. Left column: area denoted by the black dashed rectangle in d. Right column: close-ups of main lattice and superstructure peaks (white solid rectangles in the bottom-left panel), normalized to individual peak heights. Main lattice peaks stem from the undistorted lattice, whereas $\alpha$ and $\beta$ peaks are associated with the CDW-coupled periodic lattice distortion. f, Lineouts along main-$\alpha$ and main-$\beta$ directions (bottom-left panel in e) illustrate that the spot profiles of the main lattice and $\alpha$ peaks remain nearly unchanged after the optical quench.
@@
Figure 2: Generation conditions for the heterochiral CDW state. a, Normalized intensity changes of $\alpha$ (blue squares) and $\beta$ (red circles) peaks as a function of the single-pulse quench fluence (sample #2). Both traces exhibit a threshold (black vertical line) and saturation behavior (red horizontal line). b, Normalized $\beta$ peak intensity after the single-pulse quench. Measurements of the 13 samples are fitted with error functions. The obtained threshold fluences vary from 0.7 mJ cm-${}^{\textrm{-}2}$ (#3a) to 2.2 mJ cm-${}^{\textrm{-}2}$ (#5), and saturation levels range from 2 % (#3c) to about 7 % (#6), with two samples (#1 and #5) showing no signs of saturation. Considering all data, an average threshold fluence of 1.2 mJ cm-${}^{\textrm{-}2}$ and saturation intensity ratio $I_{\beta}/I_{\alpha}\approx 4\,\%$ are determined. c, Normalized intensity changes as a function of the annealing pulse fluence (sample #3a). The $\alpha$ trace does not fully return to 0 % for annealing fluences up to 1 mJ cm-${}^{\textrm{-}2}$, possibly due to some irreversible changes of the sample. d, Double-pulse optical quench (sample #4). Normalized intensity changes as a function of the pulse separation $\tau$ show an enhancement for $|\tau|\leq 400\,\textrm{fs}$ (shaded area). At a combined pulse fluence of 1.14 mJ cm-${}^{\textrm{-}2}$, the $\beta$ peak intensity generated by the double-pulse quench varies between 4.9 % ($\tau=0\,\textrm{ps}$) and 1.2 % ($\tau=-2\,\textrm{ps}$). (All data obtained by averaging the intensities of 15 CDW superstructure peaks in each LEED image. Data sets #3a, #3b, #3c taken on the same sample but with slightly different surface conditions and laser pulse durations.)
Figure 3: Proposed generation mechanism and possible structural textures for coexistent CDW chiralities. a, Suggested mechanism for the generation of a heterochiral hidden state with translational domains. The illustration depicts simplified free-energy surfaces as a function of the order parameter $\varphi$ of the homogeneous monochiral C phase, and along other coordinates of the high-dimensional configuration space (schematically indicated as multiple axes). The optically-induced displacive excitation of the CDW causes a rapid evolution of the system towards an inverted CDW state which is energetically highly unfavourable. Collision with this steep barrier leads to a forced symmetry breaking into a distribution of translational and chiral domains. Concurrent with electronic cooling and structural relaxation, this branching finally results in the formation of a metastable state with quenched disorder. b, From the relative diffraction intensities of $\alpha$ and $\beta$ peaks, the experiments imply either (Left) a nearly complete subsurface $\beta$ layer (red) underneath a surface $\alpha$ layer (blue), or (Right) a random distribution of both chiralities in the surface layer. Recent STM measurements appear to favor the latter scenario Ravnik2023 . Layers in both textures may be covered with translational domains, for which measured diffraction peak widths yield a minimum structural correlation length of about 20 nm. This implies extended interfaces between CDWs of opposite chirality at or near the surface. The emergent CDW moiré interface gives rise to a double-ring charge texture (top inset).
Figure 4: Emergent kagome system in the CDW moiré superstructure. a, Construction of the CDW moiré unit cell (see also depiction in Fig. S6b). (Left) The structure considered consists of two 1T-TaS2 trilayers of opposite CDW chirality, $\alpha$ (blue) and $\beta$ (red). (Right) The moiré structure is described by a super-hexagram unit cell containing 13 Ta hexagrams in each layer. Circles denoted by capital letters A to M indicate the center of the Ta hexagrams and the local stacking order of both layers Lee2019 , with darker filling color representing higher electrostatic energy. The inner and outer rings of the super-hexagram are marked by orange and purple, respectively. b, Kagome structural order emerging from the coupling of both layers. A comparison between atomic positions in the CDW moiré superstructure and the C-phase shows that displacements of the Ta hexagram centers (depicted by arrows) break the translational symmetry of Ta hexagram, resulting in a $13\times 13$ kagome superlattice. The concentric rings mark the inner and outer rings of the super-hexagrams. The Ta atoms at the center of the super-hexagrams experience no displacement. c, Energy- and band-resolved electron density distributions. The total electron density for bands from -$90$ meV to $60$ meV around $E_{\textrm{F}}$ develops a double-ring texture (left, red shading), while the kagome electronic subsystem (right, blue shading) features a conducting network (yellow arrows). Both charge orders follow the same kagome superlattice (black lines) as the atomic structure. Electron density of $\alpha$ and $\beta$ layers are shown in blue and red, respectively. Isosurface values are adjusted for optimal clarity of the charge pattern. d, Total density of states (DoS) and e, band structure of the heterochiral CDW state. Several flat bands and a kagome electronic subsystem (highlighted in blue) are visible near $E_{\textrm{F}}$.
Results
@@
We probe the structural changes of optically quenched 1T-TaS2 crystals with high-coherence microbeam LEED Vogelgesang2018 . Our most important observation is the emergence of CDW diffraction peaks with a chirality opposite to that of the initial state (Fig. 1d, sample #1). Specifically, we find new CDW superstructure peaks at the opposite rotation angle with respect to the main lattice after the sample is quenched by a single laser pulse. In the absence of further perturbations, this new structural feature persists beyond several hours. We denote the CDW superstructure peaks in the initial C-phase as $\alpha$ and the emergent peaks as $\beta$. The intensity difference of images before and after the optical quench (bottom-right panel of Fig. 1d) highlights the suppression of the $\alpha$ peaks and the emergence of the $\beta$ peaks in the heterochiral CDW state. A reversal to the initial LEED pattern, i.e., an erasing of the emergent $\beta$ peaks, can be achieved by illuminating the sample with a femtosecond laser pulse train for a duration of 1 s at a fluence of 0.5 mJ cm-${}^{\textrm{-}2}$ and a repetition rate of 100 kHz. Besides an occasional intensity decrease of the main lattice and $\alpha$ peaks that was observed in some samples, the resulting monochiral state has a diffraction pattern (bottom-left panel of Fig. 1d) practically indistinguishable from that of the initial C-phase.
Close-ups of the diffraction patterns before and after the quench are displayed in Fig. 1e. It is evident that the spot profiles of the main lattice and $\alpha$ peaks remain nearly unchanged after the optical quench, while the spot size of the new $\beta$ peaks is comparable to that of the $\alpha$ peaks, as shown by the lineouts in Fig. 1f. Specifically, the minimum correlation lengths of the $\alpha$ and $\beta$ chirality are determined as 22 nm and 24 nm, respectively (cf. Sec. I of Supplementary Information). Beyond the emergence of the $\beta$ chirality, we find no pronounced changes in the diffraction pattern besides a small intensity drop of the main lattice and $\alpha$ peaks. Notably, there is no evidence of substantial linewidth broadening for the main lattice or $\alpha$ peaks, which would indicate a significant increase of structural disorder. Together, this implies that the emergent $\beta$ chirality exhibits long-range structural order.
@@
Figure 2a shows the emergence of $\beta$ peaks and a concurrent suppression of the $\alpha$ peak intensity as a function of the single-pulse quench fluence. This measurement is carried out for one of the investigated samples with a randomized sequence (sample #2), demonstrating that the observations are not affected by sample deterioration. The intensity changes are normalized to the $\alpha$ peak intensity in the heterochiral state, accounting also for the background. Overall, the intensity changes are on the order of few percent of the $\alpha$ peak intensity. We find a threshold behaviour of the arising $\beta$ peak intensity at a fluence of $F_{\text{th,G}}=0.9$ mJ cm-${}^{\textrm{-}2}$, which is in close correspondence to the threshold reported for preparing the hidden state Stojchevska2014 ; Gerasimenko2019 ; Stahl2020 . Above the fluence threshold, we observe a saturation in the $\beta$ peak intensity for this sample. Neither additional features in the diffraction pattern nor permanent sample damage are observed for the quench pulse fluence in the saturation regime (no sample damage was observed in LEED images for quench pulse fluences up to 3.4 mJ cm-${}^{\textrm{-}2}$). Fluence-dependence measurements of 12 other samples (with a measurement sequence from low to high fluences) show a very similar behavior, albeit with somewhat varying threshold and saturation characteristics (Fig. 2b). We note that a comparable number of investigated samples did not show the formation of the heterochiral state, and some only did for a limited time after cleaving. Overall, a very high surface quality, low contamination, and likely low defect density appear to be a prerequisite for generating this state.
@@
By repeatedly preparing the sample in the heterochiral state using a single-pulse quench at 1 mJ cm-${}^{\textrm{-}2}$, we find a threshold for the reversal to the monochiral state at a fluence of $F_{\text{th,R}}=0.16$ mJ cm-${}^{\textrm{-}2}$ (Fig. 2c, sample #3a). In annealing the sample, a 100 kHz laser pulse train is applied for 1 s. Both the long lifetime and the comparatively low annealing threshold are clear evidence of metastability, as previously reported for the hidden state Stojchevska2014 ; Vaskivskyi2015 . The state’s high sensitivity to the surface quality also suggests that it resides in a particularly shallow metastable free-energy minimum.
@@
To elucidate the generation mechanism of the heterochiral state, we employ a double-pulse quench scheme. In Fig. 2d, the intensity changes of the $\alpha$ and $\beta$ peaks are measured (with a randomized sequence) as a function of the time delay $\tau$ between two successive quench pulses. The two pulses are orthogonally polarized, and each pulse has a below-threshold fluence of 0.57 mJ cm-${}^{\textrm{-}2}$ (sample #4, threshold fluence $F_{\text{th,G}}=1.4$ mJ cm-${}^{\textrm{-}2}$). An ultrafast enhancement of the arising $\beta$ peak intensity is found at the pulse overlap with an exponential decay time of about 400 fs, confirming the electronic origin of switching to the heterochiral state. Although the signal-to-noise ratio of our data does not allow for a determination of possible coherent oscillations, and shorter optical pulses may be necessary to resolve the CDW amplitude mode, the timescale of this enhancement is in good agreement with previous time-resolved measurements of hidden-state generation Maklar2023 ; Ravnik2018 .

### v2

# Optical switching of ferro-rotational charge-density wave states

@@
Abstract
Tailored optical excitations can steer a system along non-equilibrium pathways to metastable states with specific structural or electronic properties. The light-induced hidden state of 1T-TaS2, with its strongly enhanced conductivity and exceptionally long lifetime, represents a unique model system for studying the ultrafast switching of correlated electronic states. We use surface-sensitive electron diffraction in combination with a femtosecond optical quench to reveal the coexistence of both charge-density-wave (CDW) 2D chiralities as a structural characteristic of the hidden state, corresponding to coexisting ferro-rotational CDW states. Density functional theory (DFT) simulations of interfaces between opposite CDW 2D chiralities predict a higher-level, fractal-type moir’e superstructure with a kagome band structure near the Fermi energy. More broadly, these findings suggest that heterochiral interfaces in CDW systems provide an additional structural degree of freedom, expanding the possibilities for electronic control via twist-angle engineering.
@@ Abstract
Ultrafast control of functional materials via non-equilibrium pathways promises access to tailored electronic and structural properties [1, 2, 3, 4, 5, 6]. A particularly prominent example is the 1T-TaS2 hidden state [4], which may enable novel non-volatile all-electronic memory devices [7, 8]. The long-lived metastable hidden state exhibits metallic conductivity [4, 9, 10] and is generated from the insulating low-temperature commensurate charge-density wave phase (C-phase) [1] by applying single laser or voltage pulses [4, 14, 13, 12, 15]. Structural changes associated with this state have been probed by scanning tunneling microscopy (STM) and X-ray diffraction, yielding an emergent mosaic network of domain walls and the breakdown of interlayer dimerization [14, 13, 12, 16]. However, due to the large variability in generation conditions and properties associated with the hidden state, further systematic structural studies of metastable and mixed-order states in this material are needed, including thermally quenched pathways that stabilize a mixed-CDW regime at equilibrium [26].
@@
In the present work, we reveal a new structural characteristic concurring with the hidden state: coexistent charge-density-wave (CDW) 2D-chiralities that represent opposite domain states of a ferro-rotational order parameter defined by the in-plane orientation of the CDW superlattice. Using low-energy electron diffraction (LEED) [17] following a single-pulse optical quench [4, 15, 9, 14, 16], our experiments demonstrate the emergence of CDW domains of opposite 2D-chirality with respect to the initial state. The optically prepared state represents a multidomain configuration with imbalanced populations of two ferro-rotational domains. The electronic origin of this structural transformation is verified by double-pulse excitation [24, 18]. Fluence-dependent diffraction intensities of the newly formed domains of opposite 2D-chirality show a threshold behavior for 13 measured samples. The measured threshold fluence varies among the ensemble of crystals and is in general accordance with previously observed conditions for hidden-state generation [4, 14, 16]. At high fluence, most specimens exhibit a saturated diffraction intensity of the minority 2D-chirality between 2 % and 7 % of the majority 2D-chirality, while two samples reach about 10 % without a clear sign of saturation. Our study implies extended CDW moiré interfaces present in the laser-induced state, for which DFT simulations predict a new type of kagome charge order that supports in-plane metallic conductivity. Alongside disordered CDW stacking arrangements [13, 12], such 2D-heterochiral interfaces may contribute to the electronic properties of the hidden state, and more generally, suggest active laser control of interface 2D-chirality also in other contexts [21, 22, 19, 20].
@@
Figure 1: LEED images of the laser-induced 2D-heterochiral CDW state. a, Sketch of the optical excitation sequence, with quadrants representing the LEED images in d. b, Experimental setup for single- and double-pulse quenches monitored by LEED. c, The two 2D-chiralities of the 1T-TaS2 CDW superstructure, rotated by $\pm 13.9^{\circ}$ relative to the host lattice (black line). d, Sections of LEED images in the sequence from the C-phase to the 2D-heterochiral CDW state and back to a 2D-monochiral state. A single-pulse quench of the C-phase (top-left panel, sample #1, pulse fluence: 1.7 mJ cm-$^{\textrm{-}2}$) generates first- and second-order $\beta$ superstructure peaks (solid and dashed circles, respectively, in the top-right panel). Annealing with a laser pulse train reverses the sample to a 2D-monochiral state (bottom-left panel) indistinguishable from the initial C-phase. A difference image (bottom-right panel) highlights the emergence of $\beta$ peaks and the suppression of $\alpha$ peaks in the 2D-heterochiral state. e, Close-up LEED images of the C-phase and the 2D-heterochiral state. Left column: area denoted by the black dashed rectangle in d. Right column: close-ups of main lattice and superstructure peaks (white solid rectangles in the bottom-left panel), normalized to individual peak heights. Main lattice peaks stem from the undistorted lattice, whereas $\alpha$ and $\beta$ peaks are associated with the CDW-coupled periodic lattice distortion. f, Lineouts along main-$\alpha$ and main-$\beta$ directions (bottom-left panel in e) illustrate that the spot profiles of the main lattice and $\alpha$ peaks remain nearly unchanged after the optical quench.
@@
Figure 2: Generation conditions for the 2D-heterochiral CDW state. a, Normalized intensity changes of $\alpha$ (blue squares) and $\beta$ (red circles) peaks as a function of the single-pulse quench fluence (sample #2). Both traces exhibit a threshold (black vertical line) and saturation behavior (red horizontal line). b, Normalized $\beta$ peak intensity after the single-pulse quench. Measurements of the 13 samples are fitted with error functions. The obtained threshold fluences vary from 0.7 mJ cm-$^{\textrm{-}2}$ (#3a) to 2.2 mJ cm-$^{\textrm{-}2}$ (#5), and saturation levels range from 2 % (#3c) to about 7 % (#6), with two samples (#1 and #5) showing no signs of saturation. Considering all data, an average threshold fluence of 1.2 mJ cm-$^{\textrm{-}2}$ and saturation intensity ratio $I_{\beta}/I_{\alpha}\approx 4\,\%$ are determined. c, Normalized intensity changes as a function of the annealing pulse fluence (sample #3a). The $\alpha$ trace does not fully return to 0 % for annealing fluences up to 1 mJ cm-$^{\textrm{-}2}$, possibly due to some irreversible changes of the sample. d, Double-pulse optical quench (sample #4). Normalized intensity changes as a function of the pulse separation $\tau$ show an enhancement for $|\tau|\leq 400\,\textrm{fs}$ (shaded area). At a combined pulse fluence of 1.14 mJ cm-$^{\textrm{-}2}$, the $\beta$ peak intensity generated by the double-pulse quench varies between 4.9 % ($\tau=0\,\textrm{ps}$) and 1.2 % ($\tau=-2\,\textrm{ps}$). (All data obtained by averaging the intensities of 15 CDW superstructure peaks in each LEED image. Data sets #3a, #3b, #3c taken on the same sample but with slightly different surface conditions and laser pulse durations.)
Experimental results
@@
We probe the structural changes of optically quenched 1T-TaS2 crystals with high-coherence microbeam LEED [17]. Our most important observation is the emergence of additional CDW diffraction peaks with a 2D-chirality opposite to that of the initial state (Fig. 1d, sample #1). Specifically, we find new CDW superstructure peaks at the opposite rotation angle with respect to the main lattice after the sample is quenched by a single laser pulse. In the absence of further perturbations, this new structural feature persists beyond several hours. We denote the CDW superstructure peaks in the initial C-phase as $\alpha$ and the emergent peaks as $\beta$. These peaks correspond to the two ferro-rotational domain states of the commensurate CDW. The intensity difference of images before and after the optical quench (bottom-right panel of Fig. 1d) highlights the suppression of the $\alpha$ peaks and the emergence of the $\beta$ peaks in the 2D-heterochiral CDW state. A reversal to the initial LEED pattern, i.e., an erasing of the emergent $\beta$ peaks, can be achieved by illuminating the sample with a femtosecond laser pulse train (repetition rate 100 kHz, duration of 1 s) at a fluence of 0.5 mJ cm-$^{\textrm{-}2}$. Besides an occasional intensity decrease of the main lattice and $\alpha$ peaks that was observed in some samples, the resulting 2D-monochiral state has a diffraction pattern (bottom-left panel of Fig. 1d) practically indistinguishable from that of the initial C-phase.
As shown in Fig. 1e, the diffraction pattern after the quench reveals the emergence of $\beta$ peaks, whose spot sizes match those of the pre-existing $\alpha$ peaks (Fig. 1f). The main lattice and $\alpha$ peaks retain their profiles, with only a slight intensity decrease and no noticeable pulse-induced broadening, evidencing the absence of increased structural disorder on a scale below the coherence length of the probe [23]. The minimum correlation lengths for both $\alpha$ and $\beta$ 2D-chiralities are approximately 23 nm (see Supplementary Information, Sec. I), suggesting that the $\beta$ domains exhibit structural order beyond several CDW unit cells.
@@
Figure 2a shows the emergence of $\beta$ peaks and a concurrent suppression of the $\alpha$ peak intensity as a function of the single-pulse quench fluence. This measurement is carried out for one of the investigated samples with a randomized sequence (sample #2), demonstrating that the observations are not affected by sample deterioration. The intensity changes are normalized to the $\alpha$ peak intensity in the 2D-heterochiral state, accounting also for the background. Overall, the intensity changes are on the order of a few percent of the $\alpha$ peak intensity. We find a threshold behavior of the arising $\beta$ peak intensity at a fluence of $F_{\text{th,G}}=0.9$ mJ cm-$^{\textrm{-}2}$, which is in close correspondence to the threshold reported for preparing the hidden state [4, 14, 16]. Above the fluence threshold, we observe a saturation in the $\beta$ peak intensity for this sample. Neither additional features in the diffraction pattern nor permanent sample damage are observed for the quench fluence in the saturation regime (no sample damage was observed in LEED images for fluences up to 3.4 mJ cm-$^{\textrm{-}2}$). Fluence-dependence measurements of 12 other samples (with a measurement sequence from low to high fluences) show a very similar behavior, albeit with varying threshold and saturation characteristics (Fig. 2b). We note that a comparable number of investigated samples did not show the formation of the 2D-heterochiral state, and some only did for a limited time after cleaving. Overall, a very high surface quality, low contamination, and likely low defect density appear to be a prerequisite for generating this state.
@@
By repeatedly preparing the sample in the 2D-heterochiral state using a single-pulse quench at 1 mJ cm-$^{\textrm{-}2}$, we find an annealing threshold for the reversal to the 2D-monochiral state for a laser pulse train (repetition rate 100 kHz, duration 1 s) at a fluence of $F_{\text{th,R}}=0.16$ mJ cm-$^{\textrm{-}2}$ (Fig. 2c, sample #3a). Both the long lifetime and the comparatively low annealing threshold are clear evidence of metastability, as previously reported for the hidden state [4, 15]. The state’s high sensitivity to the surface quality also suggests that it resides in a particularly shallow metastable free-energy minimum. The moderately larger width of the CDW peaks compared to the main peaks (see Fig. S1) implies a shorter correlation length of the superstructure compared to the host lattice, even after this laser reversal to the 2D-monochiral state. This suggests the presence of residual and more stable translational $\alpha$-domains.
@@
To elucidate the generation mechanism of the 2D-heterochiral state, we employ a double-pulse quench scheme. In Fig. 2d, the intensity changes of the $\alpha$ and $\beta$ peaks are measured (with a randomized sequence) as a function of the time delay $\tau$ between two successive quench pulses. The two pulses are orthogonally polarized, and each pulse has a below-threshold fluence of 0.57 mJ cm-$^{\textrm{-}2}$ (sample #4, threshold fluence $F_{\text{th,G}}=1.4$ mJ cm-$^{\textrm{-}2}$). An ultrafast enhancement of the arising $\beta$ peak intensity is found at the pulse overlap with an exponential decay time of about 400 fs, evidencing that switching to the 2D-heterochiral state is enhanced on timescales shorter than electronic relaxation. Although the signal-to-noise ratio of our data does not allow for a determination of possible coherent oscillations, and shorter optical pulses may be necessary to resolve the CDW amplitude mode, the timescale of this enhancement is in good agreement with previous time-resolved measurements of hidden-state generation [24, 18].